# Influence of the imposed flow rate boundary condition on the flow of Bingham fluid in porous media

### 27 Dec 2023 v1

@@ I Introduction
Given the diverse applications, the flow of yield stress fluids has been the focus of numerous investigations Entov (1967); Pascal (1981); Al-Fariss T. (1985); Chase and Dachavijit (2005); Chen et al. (2005); Sochi and Blunt (2008); Chevalier et al. (2013); Talon and Bauer (2013); de Castro et al. (2014); Nash and Rees (2016); Shahsavari and McKinley (2016). A primary objective is to generalize for yield stress fluid the Darcy law that relate the total flow rate to the applied pressure drop.
From a practical point of view, there are several methods for driving flow in porous media. When the flow is through a core sample driven by pumps, the flow rate is the natural control variable. However, when the flow is driven by a height difference between fluid reservoir and inlet, the pressure difference is the control parameter Moura et al. (2020). For Newtonian flow, it is usually assumed that the classical Darcy law is independent of the prescribed boundary condition (BC) at the inlet and outlet (imposing uniform pressure or velocity). The reason behind this is that changing the boundary condition only affects the flow in a very short distance, e.g., over a few pores. It follows that if the total volume of interest is large enough, the influence of this boundary layer is negligible. The objective of this paper is to investigate the influence of the boundary condition for yield stress fluids.
@@
By prescribing the pressure difference $\Delta P$, previous studies Roux and Herrmann (1987); Chevalier and Talon (2015); Bauer et al. (2019); Liu et al. (2019) have demonstrated that the total flow rate exhibits three successive power-law regimes above a critical pressure $\Delta P_{c}$: $Q\propto(\Delta P-\Delta P_{c})^{\beta}$, where $\beta=1$ when the pressure is either close and very far from $\Delta P_{c}$. In the intermediate regime, the exponent $\beta$ is close to two Liu et al. (2019). The physical origin of this quadratic exponent lies in a progressively increasing number of channels in which there is flow as the pressure is increased. Another important point for later is the distribution of lengths of the new flow path, where longer channels are more likely to flow than shorter ones Liu et al. (2019); Schimmenti et al. (2023).
All these computational and theoretical studies were conducted with prescribed pressure at the boundaries. The aim of this research is to study the flow of yield stress fluids in porous media while imposing a prescribed flow rate. More specifically, a mixed boundary condition will be used: a uniform velocity is imposed at the inlet, while a uniform pressure is applied at the outlet. The corresponds to the experimental situation where the flow into the core sample is controlled by the rate of the pumps, whereas the other end of the core sample is open and hence kept at atmospheric pressure.
@@ II Numerical method
The solution of the system of nonlinear equations is a set of flow rates $\{q_{ij}\}$. The solution can be determined using a variational approach that minimize a functional. In this case it writes as
@@
where $\Omega$ is the ensemble of the admissible solutions satisfying conservation of mass and the prescribed flow at the inlet. Our numerical results are obtained using a modified version of this variational approach, called augmented Lagrangian method and described in Talon and Hansen Talon and Hansen (2020). Eq. (3) allows one also to obtain an important insight of the solution when $Q\rightarrow 0$ or $Q\rightarrow\infty$.
@@ III.2 Flow channels
When the flow rate is extremely low (left panel), due to the flow imposed BC, we observe a flow path starting from each inlet node. This is in contrast with the pressure imposed BC, where only a single flow path is observed. As shown in Fig. 3, the paths eventually merge with others as they progress through the system. The key feature of the low flow rates is that these paths can merge but never split. If the system is long enough, namely if $L\gg W$, the flow paths eventually converge into a single channel, as for the imposed pressure difference BC.
@@
To quantify these trends, we plot in Fig. 4 (Left) the number of flowing channels $N_{\text{channel}}$ as function of the distance, $x$, from the inlet. At low $Q$ the number of channels decreases with distance as a power power law $x^{-\delta}$, with $\delta\simeq 0.63\pm 0.05$.
@@
As the flow rate is increased, the power law decay holds within a boundary layer of length $\ell_{\text{BL}}(Q)$. Above this length, $N_{\text{channel}}$ reaches a plateau that grows as $Q^{\mu}$. Note that the existence of finite plateau is a necessary condition to enable a homogeneous description of the flow. Also, a Family-Vicsek scaling Family and Vicsek (1985) holds for $N_{\text{channel}}(x,Q)$ (see figure 4)
@@ IV Onset of the Flow: mapping to the directed polymer problem
  • For $Q\rightarrow\infty$, the flow field converges to the solution $\{q\}=\arg\min_{\{q\}\in\Omega}\sum(\frac{1}{2}\rho_{ij}q_{ij}^{2}),$, equivalent to the Newtonian flow. Hence, the mobility $\kappa_{\infty}$ corresponds to the Newtonian mobility, namely the maximal admissible mobility:
@@ V Critical exponents and non-flowing regions
In the previous section we identify the flowing channels at the onset of the flow with the ground states directed polymers in the random (porous) medium. This mapping allows us to use an efficient algorithm for the flowing channels, but it also allow to identify the critical exponent $\delta$ previously introduced with the roughness exponent $\zeta=2/3$ of the directed polymer. To do this we first characterize the statistics of the non-flowing regions using a scaling relation originally introduced for the Oslo sandpile model Paczuski and Boettcher (1996). This argument is applicable to avalanches dynamic of elastic interfaces in random media pulled at one end Aragón et al. (2016).
@@
Here $\tau_{S},\tau_{\ell}$ are the avalanche exponents with a value in the interval $(1,2)$. Both distribution have finite size cut-off ($S_{\max}\sim L^{1+\zeta}$, with $\zeta=2/3$, and $\ell_{\max}\sim L$), above which both $g_{1}(x)$ and $g_{2}(x)$ have a fast decay. We can determine the mean avalanche surface, $\langle S\rangle$ by
@@
It is also worth mentioning that at very low flow rate, the power-law relationship Eq. (5) indicates that the boundary layer should be infinite. However, due to the finite size of the system, a plateau is necessarily reached once $N_{channel}=1$. It follows a maximum length $\ell_{bc}^{max}$ for the boundary layer that depends on the domain width is $\ell_{bl}^{max}\propto W^{1/\delta}.$
@@ VI Discussion/Conclusion
In this paper we have analysed the influence of the type of boundary condition on the flow of Bingham fluid in porous media. In contrast to Newtonian fluids, the type of the applied boundary condition has a very strong influence on the flow, especially at low flow rates. One consequence is the difference in the minimal pressure for the onset of flow which is smaller when imposing the pressure imposed BC than when imposing the flow rate. Another consequence is that, at low velocity, the pressure imposition shows a single flowing channel, whereas the flow imposition displays a merging channel structure. There is therefore a boundary layer that spreads over a long distance. Its expansion increases according to a power law as the flow rate approaches zero, and increases with the system width. It follows for instance that the boundary layer, but also the difference between the two pressure of flow onset depend on the system aspect ratio. This point is particularly important in the context of homogenisation. Indeed, since it describe the system as homogeneous and derives relations between averaged quantities, the boundary conditions should have as little influence as possible. This appears to be extremely difficult in this problem because the boundary layer diverges at low flow rates, but also because it depends on the aspect ratio of the system under consideration.

### 18 Jan 2024 v2

@@ I Introduction
Given the diverse applications, the flow of yield stress fluids has been the focus of numerous investigations Entov (1967); Pascal (1981); Al-Fariss T. (1985); Chase and Dachavijit (2005); Chen et al. (2005); Sochi and Blunt (2008); Chevalier et al. (2013); Talon and Bauer (2013); de Castro et al. (2014); Chevalier et al. (2014); Bleyer and Coussot (2014); Nash and Rees (2016); Shahsavari and McKinley (2016). A primary objective is to generalize the Darcy law, that relate the total flow rate to the applied pressure drop, for yield stress fluid.
From a practical point of view, there are several methods for driving flow in porous media. When the flow is through a core sample driven by pumps, the flow rate is the natural control variable. However, when the flow is driven by a height difference between fluid reservoir and inlet, the pressure difference is the control parameter Moura et al. (2020). For Newtonian flow, it is usually assumed that the classical Darcy law is independent of the prescribed boundary condition (BC) at the inlet and outlet (imposing uniform pressure or velocity). The reason behind this is that changing the boundary condition only affects the flow in a very short distance, i.e., over a few pores. It follows that if the total volume of interest is large enough, the influence of this boundary layer is negligible. The objective of this paper is to investigate the influence of the boundary condition for yield stress fluids.
@@
By prescribing the pressure difference $\Delta P$, previous studies Roux and Herrmann (1987); Chevalier and Talon (2015); Bauer et al. (2019); Liu et al. (2019) have demonstrated that the total flow rate exhibits three successive power-law regimes above a critical pressure $\Delta P_{c}$: $Q\propto(\Delta P-\Delta P_{c})^{\beta}$, where $\beta=1$ when the pressure is either close and very far from $\Delta P_{c}$. In the intermediate regime, the exponent $\beta$ is close to two Liu et al. (2019). The physical origin of this quadratic exponent lies in a progressively increasing number of channels in which there is flow as the pressure is increased (see also Waisbord et al. (2019)). Another important point for later is the distribution of lengths of the new flow path, where longer channels are more likely to flow than shorter ones Liu et al. (2019); Schimmenti et al. (2023). It is also worth mentioning the NMR experimental work of Chevalier et al. Chevalier et al. (2014) where no fluid at rest was observed even at low flow rates. The authors then suggest that the breakdown of the non-Newtonian character may be due to the boundary condition.
Indeed, all these computational and theoretical studies were conducted with prescribed pressure at the boundaries. The aim of this research is thus to study the flow of yield stress fluids in porous media while imposing a prescribed flow rate. More specifically, a mixed boundary condition will be used: a uniform velocity is imposed at the inlet, while a uniform pressure is applied at the outlet. This corresponds to the experimental situation where the flow into the core sample is controlled by the rate of the pumps, whereas the other end of the core sample is open and hence kept at atmospheric pressure.
@@ II Numerical method
The solution of the system of nonlinear equations is a set of flow rates $\{q_{ij}\}$. The solution can be determined using a variational approach that minimize a functional Talon and Hansen (2020); Talon (2022). In this case it writes as
@@
where $\Omega$ is the ensemble of the admissible solutions satisfying conservation of mass and the prescribed flow at the inlet. Our numerical results are obtained using a modified version of this variational approach, called augmented Lagrangian method and is described in Talon and Hansen Talon and Hansen (2020). Eq. (3) allows one also to obtain an important insight of the solution when $Q\rightarrow 0$ or $Q\rightarrow\infty$.
@@ III.2 Flow channels
When the flow rate is extremely low (left panel), due to the flow imposed BC, we observe a flow path starting from each inlet node. This is in contrast with the pressure imposed BC, where only a single flow path is observed. As shown in Fig. 3, the paths eventually merge with others as they progress through the system. The key feature of the low flow rates is that these paths merge but never split. If the system is long enough, namely if $L\gg W$, the flow paths eventually converge into a single channel, as for the imposed pressure difference BC.
@@
To quantify these trends, we plot in Fig. 4 (Left) the number of flowing channels $N_{\text{channel}}$ as function of the distance, $x$, from the inlet. At low $Q$ the number of channels decreases with distance as a power law $x^{-\delta}$, with $\delta\simeq 0.63\pm 0.05$.
@@
As the flow rate is increased, the power law decay holds within a boundary layer of length $\ell_{\text{BL}}(Q)$. Above this length, $N_{\text{channel}}$ reaches a plateau that grows as $Q^{\mu}$. Note that the existence of a plateau is a necessary condition to enable a homogeneous description of the flow. Also, a Family-Vicsek scaling Family and Vicsek (1985) holds for $N_{\text{channel}}(x,Q)$ (see figure 4)
@@ IV Onset of the Flow: mapping to the directed polymer problem
  • For $Q\rightarrow\infty$, the flow field converges to the solution $\{q\}=\arg\min_{\{q\}\in\Omega}\sum(\frac{1}{2}\rho_{ij}q_{ij}^{2}),$ equivalent to the Newtonian flow. Hence, the mobility $\kappa_{\infty}$ corresponds to the Newtonian mobility, namely the maximal admissible mobility:
@@ V Critical exponents and non-flowing regions
In the previous section we identify the flowing channels at the onset of the flow with the ground states directed polymers in the random (porous) medium. This mapping allows us to use an efficient algorithm for the flowing channels, but it also allows us to identify the critical exponent $\delta$ previously introduced with the roughness exponent $\zeta=2/3$ of the directed polymer. To do this we first characterize the statistics of the non-flowing regions using a scaling relation originally introduced for the Oslo sandpile model Paczuski and Boettcher (1996). This argument is applicable to avalanche dynamics of elastic interfaces in random media pulled at one end Aragón et al. (2016).
@@
Here $\tau_{S},\tau_{\ell}$ are the avalanche exponents with a value in the interval $(1,2)$. Both distributions have a finite size cut-off ($S_{\max}\sim L^{1+\zeta}$, with $\zeta=2/3$, and $\ell_{\max}\sim L$), above which both $g_{1}(x)$ and $g_{2}(x)$ have a fast decay. We can determine the mean avalanche surface, $\langle S\rangle$ by
@@
It is also worth mentioning that at very low flow rate, the power-law relationship Eq. (5) indicates that the boundary layer should be infinite. However, due to the finite size of the system, a plateau is necessarily reached once $N_{channel}=1$. It follows a maximum length $\ell_{bc}^{max}$ for the boundary layer that depends on the domain width is $\ell_{bc}^{max}\propto W^{1/\delta}.$
@@ VI Discussion/Conclusion
In this paper we have analysed the influence of the type of boundary condition on the flow of Bingham fluid in porous media. In contrast to Newtonian fluids, the type of the applied boundary condition has a very strong influence on the flow, especially at low flow rates. One consequence is the difference in the minimal pressure for the onset of flow which is smaller when imposing the pressure imposed BC than when imposing the flow rate. Another consequence is that, at low velocity, the pressure imposition shows a single flowing channel, whereas the flow imposition displays a merging channel structure. There is therefore a boundary layer that spreads over a long distance. Its expansion increases according to a power law as the flow rate approaches zero, and increases with the system width. It follows for instance that the boundary layer, but also the difference between the two pressure of flow onset depend on the system aspect ratio. This point is particularly important in the context of homogenisation. Indeed, since it describe the system as homogeneous and derives relations between averaged quantities, the boundary conditions should have as little influence as possible. This appears to be extremely difficult in this problem because the boundary layer diverges at low flow rates, but also because it depends on the aspect ratio of the system under consideration. In the light of this study, it is difficult to claim that the boundary condition is responsible for the breakdown of the non-Newtonian behaviour observed in Chevalier et al. (2014), since we still observe non-Newtonian behaviour with some fluid at rest in the boundary layer. However, it is worth recalling that this path selection is closely related to the disorder of the system (see Chevalier and Talon (2015); Waisbord et al. (2019); Kostenko and Talon (2019)). It is then probably the combination of both the imposed flow rate condition and the use of a relatively homogeneous porous medium (monodisperse glass bead) that makes the detection of such phenomena difficult.
@@
Acknowledgements
The authors thanks F. Lanza, S. Sinha and Ph. Coussot for fruitfull discussions. This work was partly supported by the Research Council of Norway through its Center of Excellence funding scheme, project number 262644. Further support, also from the Research Council of Norway, was provided through its INTPART program, project number 309139.